Victory for MSU! ~ Sonya believes that a larger proportion of undergraduate students know the MSU fight song by heart than do graduate students. She randomly samples 201 undergraduate students and 98 graduate students and finds that 145 undergraduate students and 60 graduate students know the MSU fight song by heart.

Sonya wants to determine if the proportion of undergraduate students who know the fight song is larger than the proportion of graduate students.

Notation: 1=undergraduate and 2=graduate.

A normal distribution is used to carry out a hypothesis test for this scenario.

What is the pooled estimate for the common value of the population proportion?

Your Answer:
Answer #1

For undergraduate students: n1 = 201; x1 = 145

For graduate students: n2 = 98; x2 = 60

The pooled estimate for the common value of the population proportion

= (x1+x2)/(n1+n2)

= (145+60)/(201+98)

= 205/299

= 0.6856187 (Rounded as needed)